Mardi received an inheritance of $50,000. She invested part at 7% and deposited the remainder in tax-free bonds at 12%. Her total annual income from the investments was $4500 Find the amount invested at 7%.
Answer by palanisamy(496)

Mardi received an inheritance of $50,000
Let the amount invested at 7% be = x
She,deposited the remainder in tax-free bonds at 12%
So the amount invested at 12% in tax-free bonds = 50000-x
Total annual income is x*(7/100)+(50000-x)*(12/100) = 4500
7x+600000-12x = 450000
-5x = 450000-600000
-5x = -150000
x = -150000/(-5)
x = 30000
So the amount invested at 7% = $30,000